OpenStudy (anonymous):

After picking apples, 4 friends compared the amount of apples in their buckets. Dan’s bucket was 2/3 full, Mary’s bucket was 4/5 full, Julie’s bucket was 3/4 full, and Jane’s bucket was 1/8 full. How would these amounts be arranged in order from greatest to least?

OpenStudy (anonymous):

2/3=.66667 4/5=.8 3/5=.75 1/8=.125 4/5,3/5,2/3,1/8